# spacing=nonfrench

### 04 Mar 2026 v1

@@ 2.5.2. Transport of structure
In order to define a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ transporting $f$ to $f_{H}$, we begin by matching $S_{f}$ with $S_{H}$. Set $j\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=p^{\ell}-p^{\ell-1}$ and choose a system $(\alpha_{1},\dots,\alpha_{j})$ of elements of $S_{f}$ inducing a partition of $S_{f}$ as in §2.1. Similarly, pick a system $(\beta_{1},\dots,\beta_{j})$ of elements of $S_{H}$ where we run through the torsion points of $H$ of exact order $p^{\ell}$. Such a system, in turn, induces a partition of $S_{H}$. For each integers $1\leqslant i\leqslant j$ and $n\geqslant 0$, and all $g\in G_{L}$, set
(3) $\tau((f^{\circ n}\circ[\chi_{f,\alpha_{i}}(g)]_{f})(\alpha_{i}))\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(f_{H}^{\circ n}\circ[\chi_{f}(g)]_{H})(\beta_{i})$
and $\tau(0)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=0$. Fixing $n=0$ and taking $i\in\{1,\dots,j\}$ in (3) gives a well-defined bijection $\tau\colon S_{f,\alpha_{i}}\rightarrow S_{H,\beta_{i}}$ since the $\mathbb{Z}_{p}$-iterates of $u$ act simply transitively on $S_{f,\alpha_{i}}$, and the same holds with $u_{H}$ and $S_{f,\beta_{i}}$. Varying $n\geqslant 0$ in (3) allows us to reach the other points of $T_{f}$ and $\operatorname{T}_{p}(H)$. This defines a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ associating $f$ with $f_{H}$, and the statement follows by transport of structure. ∎
Remark 2.10.
A different choice of representatives for the partitions in the previous proof would yield the same $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$, but would permute the roles of the individual elements. This is immaterial for our purpose, since what matters to us are the group law and the property that the power series $f$ remains an endomorphism of $T_{f}$ after transport of structure, which are unaffected by such choices. Nevertheless, the latent formal group $F$ obtained in Theorem 3.1 will satisfy $\operatorname{T}_{p}(F)=T_{f}$ as $G_{L}$-sets, but only $\operatorname{T}_{p}(F)\simeq T_{f}$ as $\mathbb{Z}_{p}[G_{L}]$-modules.
3. The latent formal group
Possessing a $p$-adic Tate module and having access to the existence of a quasi-inverse to the equivalence in §2.2.5 is not sufficient to solve Lubin’s conjecture: it is not enough to work up to isomorphism; one must have equality. For this purpose, each time a functor is applied to $T_{f}$, it is necessary to keep track of the behavior of $f$. Recall that $f$ acts on $T_{f}$ by componentwise evaluation, or equivalently, by multiplication by its linear coefficient, and that all its other coefficients are uniquely determined by the action of $\mathbb{Z}_{p}$ on $T_{f}$. Our strategy requires a careful analysis of the explicit constructions in integral $p$-adic Hodge theory. One such approach proceeds through the work of Kisin in [Kis06] and [Kis09] to obtain a formal Breuil-Kisin module, and then via the one of Zink in [Zin01] and [Zin02] on displays, indicated by the dashed path in the diagram:
[…]
where the categories and functors appearing therein will be defined progressively throughout this section. At each stage, we track the action of $f$ on the object being constructed and ensure that it indeed encodes the data of its other coefficients just as $T_{f}$ does. To this end, we construct inside or extend the previous object. These precautions are necessary in order to conclude that the formal group obtained in the end is the latent one.
@@ 3.1. The associated Breuil-Kisin module
Recall that we denote by $L_{0}$ the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$, and $\sigma$ its Frobenius. We have also fixed a uniformizer $\omega_{L}$ of $L$, and we let $E(X)$ denote its Eisenstein minimal polynomial over $L_{0}$. Let $L_{\infty}\!$ be the extension generated over $L$ by the components of a compatible sequence of $p$-th roots of $\omega_{L}$ in $\mathbb{C}_{p}$. We define two subrings $\mathfrak{S}\subset\mathfrak{S}^{\mathrm{ur}}$ of $\operatorname{W}(\mathcal{R})$ as follows. Let
$\mathfrak{S}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{W}(\mathbb{F}_{q})[\![X]\!]$
be the Kisin ring, endowed with a $\sigma$-semilinear Frobenius endomorphism $\varphi$ defined by $\varphi(X)=X^{p}$. According to [Kis06, §2.1.1], there exists an embedding $\mathfrak{S}\hookrightarrow\operatorname{W}(\mathcal{R})$ that preserves the Frobenius. Let $\mathscr{E}$ be the fraction field of the $p$-adic completion of $\mathfrak{S}[1/X]$. The above embedding extends to an embedding $\mathscr{E}\hookrightarrow\operatorname{W}(\operatorname{Frac}(\mathcal{R}))$, and we denote by $\mathscr{E}^{\mathrm{ur}}$ the maximal unramified extension of $\mathscr{E}$ inside $\operatorname{W}(\operatorname{Frac}(\mathcal{R}))[1/p]$, and by $\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}$ its ring of integers. We then define
$\mathfrak{S}^{\mathrm{ur}}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}\cap\operatorname{W}(\mathcal{R}).$
@@ 3.1.1. Breuil-Kisin modules
A Breuil-Kisin module $(\mathfrak{M},\varphi)$ is a finite free $\mathfrak{S}$-module $\mathfrak{M}$ equipped with a $\varphi_{\mathfrak{S}}$-semilinear Frobenius endomorphism $\varphi\colon\mathfrak{M}\rightarrow\mathfrak{M}$ whose cokernel of its $\mathfrak{S}$-linearization $1\otimes\varphi\colon\mathfrak{S}\otimes_{\mathfrak{S},\varphi}\mathfrak{M}\rightarrow\mathfrak{M}$ is killed by $E(X)$. Together with $\varphi$-equivariant $\mathfrak{S}$-linear maps, these objects form the category $\operatorname{\mathbf{BT}}_{\mathfrak{S}}(\varphi)$. Furthermore, Kisin obtained in [Kis06] a functor $\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}$ such that the following diagram commutes:
@@
where $D$ denotes the functor defined in [Kis06, §1], while $\{0,1\}$” stands for the weights, or the jumps, in $\{0,1\}$. The full subcategory of Breuil-Kisin modules corresponding to the representations of weight $1$ is denoted by $\operatorname{\mathbf{BT}}_{\mathfrak{S}}^{\mathrm{f}}(\varphi)$, and its objects are called formal Breuil-Kisin modules.
@@ 3.1.2. Construction
Set $M_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Hom}_{\mathbb{Z}_{p}[G_{L_{\infty}}]}(T_{f},\mathfrak{S}^{\mathrm{ur}})$ to be the free $\mathfrak{S}$-module of rank $1$ appearing in [Kis06, Lemma 2.1.10]. The action of the power series $f$ on $T_{f}$ extends to $M_{f}$ to be the multiplication by its linear coefficient. Indeed, given a generator $\pi_{f}\mapsto s$ of $M_{f}$, one has $f(\pi_{f}\mapsto s)=f^{\prime}(0)(\pi_{f}\mapsto s)$ because $f(\pi_{f})=f^{\prime}(0)\pi_{f}$. We do not loose any information on the remaining coefficients of $f$ because they are encoded in the action of $\mathbb{Z}_{p}\subset\mathfrak{S}$ on $M_{f}$. We let
$\mathfrak{M}_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}(T_{f})$
which is a rank-$1$ formal Breuil-Kisin module. By [Kis06, Lemma 2.1.10], we may assume that $\mathfrak{M}_{f}\subseteq M_{f}$ as $\mathfrak{S}$-modules. Consequently, the action of $f$ on $M_{f}$ restricts to $\mathfrak{M}_{f}$, and the inclusion guarantees that the latter encapsulates the other coefficients of the formal power series $f$.
@@ 3.2. The associated window
Let $S$ be the $p$-adic completion of the PD-envelope of $\operatorname{W}(\mathbb{F}_{q})[X]$ with respect to the ideal $(E(X))$, and let $\operatorname{Fil}^{1}S\subset S$ be the closure of the ideal generated by the divided powers of $E(X)$. As topological rings, we have an identification $S/\operatorname{Fil}^{1}S\simeq\mathcal{O}_{L}$. By construction, we have an inclusion with the Kisin ring $\mathfrak{S}\subset S$, and the Frobenius $\varphi$ on $\mathfrak{S}$ naturally extends to a $\sigma$-semilinear Frobenius endomorphism $\varphi$ on the ring $S$.
@@ 3.2.1. $S$-windows
Let $R\in\{\mathcal{O}_{L}/(p^{n})\,;\,n\geqslant 0\}$ be a ring. An $S$-window $(W,W^{\prime},\varphi)$ over $R$ is a finite projective $S$-module $W$ together with a distinguished submodule $W^{\prime}$ containing $(\operatorname{Fil}^{1}S)W$ such that $W/W^{\prime}$ is a projective module over $R$, and a $\sigma$-semilinear Frobenius endomorphism $\varphi\colon W\rightarrow W$ satisfying $\varphi(W^{\prime})\subseteq pW$; $W$ is generated as an $S$-module by $\varphi(W)+\varphi/p(W^{\prime})$. Given that $pW\subseteq(1\otimes\varphi)(\varphi^{*}(W))$, with $\varphi^{*}(W)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=S\otimes_{S,\varphi}W$, the map $1\otimes\varphi$ is injective and there exists a unique $S$-linear map $\psi\colon W\rightarrow\varphi^{*}(W)$ such that the composite $(1\otimes\varphi)\circ\psi\colon W\rightarrow W$ is the multiplication-by-$p$ map. We further require, for $n$ sufficiently large, that
@@
to factor through $(p,\operatorname{Fil}^{1}S)\varphi^{n*}(W)$. Such objects form the category $S\textbf{-}\mathbf{Win}/R$, whose morphisms are $\varphi$-equivariant $S$-linear maps that preserve the distinguished submodule. In the particular case where $R=\mathcal{O}_{L}$, the data of an $S$-window over $\mathcal{O}_{L}$ is equivalent to
the data of a system $(W_{n})_{n\geqslant 1}$, where $W_{n}$ is an $S$-window over $\mathcal{O}_{L}/(p^{n})$ and is obtained from $W_{n+1}$ by base change for every integer $n\geqslant 1$. The extension
@@ 3.2.2. Construction
We set
$W_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=S\otimes_{\varphi,\mathfrak{S}}\mathfrak{M}_{f}$
to be the $S$-window over $\mathcal{O}_{L}$ associated to our formal Breuil-Kisin module. Since Zink’s functors in [Zin01] for constructing $p$-divisible groups are formulated in terms of the limit description, we shall work with this perspective. Hence let $(W_{f,n})_{n\geqslant 1}$ be the corresponding system. Upon passing to the quotient of $W_{f}$ with which $W_{f,n}$ is identified, we may assume that each $W_{f,n}$ is a quotient of $W_{f}$. The action of $f$ on $\mathfrak{M}_{f}$ extends to an action of $f$ on $W_{f}$ by the multiplication-by-its-linear-coefficient map. This action descends compatibly to each quotient $W_{f,n}$. Taking the $W_{f,n}$ as quotients of $W_{f}$ guarantees that the remaining coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subset S$ on each of these objects.
@@ 3.3. The associated $p$-divisible group
Let $(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$ be the category of finite-height connected $p$-divisible groups over $\mathcal{O}_{L}$. In [Zin01], Zink introduced two functors, $\operatorname{Dsp}$ and $\operatorname{BT}$, such that
@@
is an equivalence. These functors are defined on $S$-windows over $\mathcal{O}_{L}$ expressed as a limit. Let $n\geqslant 1$ and write $R_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{O}_{L}/(p^{n})$. Set $P_{f,n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Dsp}(W_{f,n})=\operatorname{W}(R_{n})\otimes_{\varkappa_{n},S}W_{f,n}$ which is a $\operatorname{W}(R_{n})$-module, where $\varkappa_{n}\colon S\rightarrow\operatorname{W}(R_{n})$ is the map defined in [Zin01, p. 499]. The action of $f$ on $W_{f,n}$ extends to an action on $P_{f,n}$ and again is the multiplication-by-its-linear-coefficient map. All other coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subset\operatorname{W}(R_{n})$ on $P_{f,n}$. Then, we set
@@
which is a height-$1$ connected $p$-divisible group over $\mathcal{O}_{L}$. Denoting by $\operatorname{W}_{\mathrm{f}}$ the functor of Witt vectors with finite support, one defines $\Gamma_{f,n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{BT}(P_{f,n})$ as a certain quotient of the functor $\operatorname{W}_{\mathrm{f}}(-)\otimes_{\operatorname{W}(R_{n})}P_{f,n}$. The action of $f$ on $P_{f,n}$ extends to the latter functor, descends to the quotient $\Gamma_{f,n}$ and is compatible with the transition maps. Consequently, the action of $f$ passes to the limit $\Gamma_{f}$ and is the multiplication-by-its-linear-coefficient map. The remaining coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}$, that is, by the structural data of $\Gamma_{f}$ arising from the structure of $T_{f}$.

### 22 Apr 2026 v2

@@ 3. The latent formal group
where the categories and functors appearing therein will be defined progressively throughout this section. At each stage, we track the action of $f$ on the object being constructed and ensure that it indeed encodes the data of its other coefficients just as $T_{f}$ does. These precautions are necessary in order to conclude that the formal group obtained in the end is the latent one.
@@ 3.1. The associated Breuil-Kisin module
Recall that we let $L_{0}$ be the maximal unramified extension of $\mathbb{Q}_{p}$ inside $L$ and let $\sigma$ be its Frobenius. We have also fixed a uniformizer $\pi_{L}$ of $L$, and write $E(X)$ for its Eisenstein minimal polynomial over $L_{0}$. Denote by $L_{\infty}\!$ the extension generated over $L$ by the components of a compatible sequence of $p$-th roots of $\pi_{L}$ in $\mathfrak{m}_{\mathbb{C}_{p}\!}$. We define two subrings $\mathfrak{S}\subset\mathfrak{S}^{\mathrm{ur}}$ of $\operatorname{W}(\mathcal{R})$ as follows. Let
$\mathfrak{S}=\operatorname{W}(\mathbb{F}_{q})[\![X]\!]$
be the Kisin ring, endowed with a $\sigma$-semilinear Frobenius endomorphism $\varphi$ defined by $\varphi(X)=X^{p}$. According to [Kis06, §2.1.1], there exists an embedding $\mathfrak{S}\hookrightarrow\operatorname{W}(\mathcal{R})$ that preserves the Frobenius. Let $\mathscr{E}$ be the fraction field of the $p$-adic completion of $\mathfrak{S}[1/X]$. The previous embedding extends to an embedding $\mathscr{E}\hookrightarrow\operatorname{W}(\operatorname{Frac}(\mathcal{R}))$, and we denote by $\mathscr{E}^{\mathrm{ur}}$ the maximal unramified extension of $\mathscr{E}$ inside $\operatorname{W}(\operatorname{Frac}(\mathcal{R}))[1/p]$ and by $\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}$ its ring of integers. We then define
$\mathfrak{S}^{\mathrm{ur}}=\mathcal{O}_{\mathscr{E}^{\mathrm{ur}}}\cap\operatorname{W}(\mathcal{R}).$
@@ 3.1.1. Breuil-Kisin modules
A Breuil-Kisin module $(\mathfrak{M},\varphi)$ is a finite free $\mathfrak{S}$-module $\mathfrak{M}$ equipped with a $\varphi_{\mathfrak{S}}$-semilinear Frobenius endomorphism $\varphi\colon\mathfrak{M}\rightarrow\mathfrak{M}$ whose cokernel of its $\mathfrak{S}$-linearization $1\otimes\varphi\colon\mathfrak{S}\otimes_{\mathfrak{S},\varphi}\mathfrak{M}\rightarrow\mathfrak{M}$ is killed by $E(X)$. Together with $\varphi$-equivariant $\mathfrak{S}$-linear maps, these objects form the category $\operatorname{\mathbf{BT}}_{\mathfrak{S}}(\varphi)$. Kisin obtained in [Kis06] a functor $\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}$ such that the following diagram commutes:
@@
where $D$ is the functor defined in [Kis06, §1] while $\{0,1\}$” stands for the weights or the jumps in $\{0,1\}$. The full subcategory of Breuil-Kisin modules corresponding to the representations of formal group type is denoted by $\operatorname{\mathbf{BT}}_{\mathfrak{S}}^{\mathrm{f}}(\varphi)$, and its objects are called formal Breuil-Kisin modules.
@@ 3.1.2. Construction
Set
$M_{f}=\operatorname{Hom}_{\mathbb{Z}_{p}[G_{L_{\infty}}]}(T_{f},\mathfrak{S}^{\mathrm{ur}})$
to be the free $\mathfrak{S}$-module of rank $1$ appearing in [Kis06, Lemma 2.1.10]. The action of $f$ on $T_{f}$ extends to $M_{f}$ to be the multiplication by its linear coefficient. Indeed, given a generator $\pi\mapsto s$ of $M_{f}$, one has $f(\pi\mapsto s)=f^{\prime}(0)(\pi\mapsto s)$ because $f(\pi)=f^{\prime}(0)\pi$. We do not loose any information on the remaining coefficients of $f$ since they are encoded in the action of $\mathbb{Z}_{p}\subset\mathfrak{S}$ on $M_{f}$. We let
$\mathfrak{M}_{f}=\operatorname{\mathbf{D}}_{\mathfrak{S}}^{*}(T_{f})$
which is a rank-one formal Breuil-Kisin module. By [Kis06, Lemma 2.1.10], we may assume that $\mathfrak{M}_{f}\subseteq M_{f}$ as $\mathfrak{S}$-modules. Therefore the action of $f$ on $M_{f}$ restricts to $\mathfrak{M}_{f}$ and the inclusion guarantees that the latter encapsulates the other coefficients of the formal power series $f$.
@@ 3.2. The associated window
Let $S$ be the $p$-adic completion of the PD-envelope of $\operatorname{W}(\mathbb{F}_{q})[X]$ with respect to the ideal $(E(X))$, and let $\operatorname{Fil}^{1}S\subset S$ be the closure of the ideal generated by the divided powers of $E(X)$. As a topological rings, one has an identification $S/\operatorname{Fil}^{1}S\simeq\mathcal{O}_{L}$. By definition, one has $\mathfrak{S}\subset S$ and the Frobenius $\varphi$ on $\mathfrak{S}$ extends to a $\sigma$-semilinear Frobenius endomorphism $\varphi$ on the ring $S$.
@@ 3.2.1. $S$-windows
Let $R\in\{\mathcal{O}_{L}/(p^{n})\,;\,n\geqslant 0\}$ be a ring. An $S$-window $(W,W^{\prime},\varphi)$ over $R$ is a finite projective $S$-module $W$ together with a submodule $W^{\prime}$ containing $(\operatorname{Fil}^{1}S)W$ and a $\sigma$-semilinear Frobenius endomorphism $\varphi\colon W\rightarrow W$ such that:
  1. (1)
    $W/W^{\prime}$ is a projective $R$-module,
  2. (2)
    $\varphi(W^{\prime})\subseteq pW$ and $W$ is generated by $\varphi(W)+\varphi/p(W^{\prime})$ as an $S$-module,
  3. (3)
    since $pW\subseteq(1\otimes\varphi)(\varphi^{*}(W))$, where $\varphi^{*}(W)=S\otimes_{S,\varphi}W$, the map $1\otimes\varphi$ is injective and there exists a unique $S$-linear map $\psi\colon W\rightarrow\varphi^{*}(W)$ such that $(1\otimes\varphi)\circ\psi\colon W\rightarrow W$ is the multiplication-by-$p$ map. Then, for $n$ sufficiently large,
@@
    factors through $(p,\operatorname{Fil}^{1}S)\varphi^{n*}(W)$.
Together with $\varphi$-equivariant $S$-linear maps preserving the distinguished submodule, these objects form the category $S\textbf{-}\mathbf{Win}/R$. In the particular case of $R=\mathcal{O}_{L}$, the data of an $S$-window over $\mathcal{O}_{L}$ is equivalent to the data of a system $(W_{n})_{n\geqslant 1}$, where $W_{n}$ is an $S$-window over $\mathcal{O}_{L}/(p^{n})$ and is obtained from $W_{n+1}$ by base change for every integer $n\geqslant 1$. The extension
@@ 3.2.2. Construction
Set
$W_{f}=S\otimes_{\varphi,\mathfrak{S}}\mathfrak{M}_{f}$
to be the $S$-window over $\mathcal{O}_{L}$ associated to our formal Breuil-Kisin module. Since Zink’s functors in [Zin01] for constructing $p$-divisible groups are formulated in terms of the limit description, we shall work with this perspective. Hence let $(W_{f,n})_{n\geqslant 1}$ be the corresponding system. Upon passing to the quotient of $W_{f}$ with which $W_{f,n}$ is identified, we may assume that every $W_{f,n}$ is a quotient of $W_{f}$. The action of $f$ on $\mathfrak{M}_{f}$ extends to an action of $f$ on $W_{f}$ to be the multiplication-by-$f^{\prime}(0)$ map. This action descends compatibly to each quotient $W_{f,n}$. Taking the $W_{f,n}$ as quotients of $W_{f}$ guarantees that the remaining coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subset S$ on each of these objects.
@@ 3.3. The associated $p$-divisible group
Denote by $(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$ the category of finite-height connected $p$-divisible groups over $\mathcal{O}_{L}$. In [Zin01], Zink introduced two functors, namely $\operatorname{Dsp}$ and $\operatorname{BT}$, such that
@@
is an equivalence. These functors are defined on $S$-windows over $\mathcal{O}_{L}$ expressed as a limit. Let $n\geqslant 1$ and $R_{n}=\mathcal{O}_{L}/(p^{n})$. Set $P_{f,n}=\operatorname{Dsp}(W_{f,n})=\operatorname{W}(R_{n})\otimes_{\varkappa_{n},S}W_{f,n}$ which is a $\operatorname{W}(R_{n})$-module, with $\varkappa_{n}\colon S\rightarrow\operatorname{W}(R_{n})$ the map of [Zin01, p. 499]. The action of $f$ on $W_{f,n}$ extends to an action on $P_{f,n}$ and, again, is the multiplication-by-its-linear-coefficient map. All the other coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}\subseteq\operatorname{W}(R_{n})$ on $P_{f,n}$. Then we set
@@
which is a height-one connected $p$-divisible group over $\mathcal{O}_{L}$. Let $\operatorname{W}_{\mathrm{f}}$ be the functor of Witt vectors with finite support, one defines $\Gamma_{f,n}=\operatorname{BT}(P_{f,n})$ as a certain quotient of the functor $\operatorname{W}_{\mathrm{f}}(-)\otimes_{\operatorname{W}(R_{n})}P_{f,n}$. The action of $f$ on $P_{f,n}$ extends to the previous functor, descends to the quotient $\Gamma_{f,n}$ and is compatible with the transition maps. Consequently the action of $f$ passes to the limit $\Gamma_{f}$ and is the multiplication-by-$f^{\prime}(0)$ map. The other coefficients of $f$ are uniquely determined by the action of $\mathbb{Z}_{p}$, that is, by the structural data of $\Gamma_{f}$ arising from the structure of $T_{f}$.